If a single die is rolled (a cube where the sides are numbered 1 through 6), the probability of rolling at least a 3 is 0.33.
Answer by stanbon(75887)

If a single die is rolled (a cube where the sides are numbered 1 through 6), the probability of rolling at least a 3 is 0.33.
-----
# of ways to succeed by rolling 3,4,5 or 6::: 4
------
# of possible outcomes: 6
----
P(roll at least a 3) = 4/6 = 2/3 = 0.666666...
Cheers,
Stan H.
============